Riverside Inc. makes one model of wooden canoe. Partial information for it follows: Number of Canoes Produced and Sold 500 65Required 1 Required 3 Required 4 Suppose Riverside sells its canoes for $508 each. Calculate the contribution margin per canoComplete this question by entering your answers in the tabs below. Required 1 Required 3 Required 4 Complete the table. (RounRiverside Inc. makes one model of wooden canoe. Partial information for it follows: Number of Canoes Produced and Sold 500 65Calculations would be greatly appreciated as well!
Answer #1

1) Table

500 650 800
Total Cost
Variable cost 67500 87750 108000
Fixed cost 149500 149500 149500
Total Cost 217000 237250 257500
Cost per unit
Variable cost 135 135 135
Fixed cost 299 230 186.88
Total Cost per unit 434 365 321.88

b) Unit Contribution margin = 508-135 = 373

Contribution margin ratio = 373/508 = 73%

c) Contribution margin income statement

Sales (850*508) 431800
Variable cost (850*135) 114750
Contribution margin 317050
Fixed cost 149500
Operating income 167550